OpenStudy (anonymous):

Nathan has 80 stamps in his collection. He adds 1 stamp to it today. Each day he plans to add twice the number of stamps as the previous day. If he keeps adding stamps at this rate for n days, which recursive function represents the number of stamps he has on any day in the future? The recursive function is_, starting at_ . A) next=now+2^n A)80 B) next=now+2^n-1 B) 2^n C) next=now-2^n-1 C)80^n

OpenStudy (anonymous):

help please

OpenStudy (theloshua):

I believe A

OpenStudy (anonymous):

for which column?? @Theloshua

OpenStudy (theloshua):

The left one

OpenStudy (kropot72):

If A in the left column is chosen, then after 1 day he will have 82 stamps. 80 + 2^(1) = 82. But he should have 81 stamps after 1 day.

OpenStudy (kropot72):

So which of the choices will give 81 stamps after one day?

OpenStudy (anonymous):

C??? @kropot72

OpenStudy (kropot72):

C will give 79 stamps after 1 day. 80 - 2^(n-1)=80 - 1 = 79